The larger of the two angles made with x-axis of a straight line drawn through (1,2) so that it intersects x+y= 4 at a distance (√6 ) / 3 from (1,2) is : -

Grade:11

1 Answers

Arun
6 years ago
suppose a line passing through (1,2)makes an angle A with x-axis.then its eqn is
x-1/cosA =y-2/sinA
the coordinates of pt on line at a distance
6/3=√2/3 from (1,2)are given by
x-1/cosA =y-2/sinA=√2/3 or ,(1+√2/3*cosA ,2+
2/3*sinA)
this pt lies on x+y =4
cosA+sinA=√3/2
sin(A+\huge \pi/4)=√3/2
A+\huge \pi/4=\huge \pi/3,2\huge \pi/3
A=\huge \pi/12,5\huge \pi/12
A=15 or 75